OpenStudy (anonymous):

Related rates problem: If a snowball melts so that its surface area decreases at a rate of 1cm^2/min, find the rate at which the diameter decreases when the diameter is 10cm.

OpenStudy (phi):

how about this let D= diameter (cm), S= surface area, t= time (minutes) \[\frac{dD}{dt}= \frac{dD}{dS}\frac{dS}{dt}\] use \[ S= \pi D^2 \]